# Visibility-Aware Keypoint Localization for 6DoF Object Pose Estimation

### v1

@@ Abstract
Localizing predefined 3D keypoints in a 2D image is an effective way to establish 3D-2D correspondences for 6DoF object pose estimation. However, unreliable localization results of invisible keypoints degrade the quality of correspondences. In this paper, we address this issue by localizing the important keypoints in terms of visibility. Since keypoint visibility information is currently missing in dataset collection process, we propose an efficient way to generate binary visibility labels from available object-level annotations, for keypoints of both asymmetric objects and symmetric objects. We further derive real-valued visibility-aware importance from binary labels based on PageRank algorithm. Taking advantage of the flexibility of our visibility-aware importance, we construct VAPO (Visibility-Aware POse estimator) by integrating the visibility-aware importance with a state-of-the-art pose estimation algorithm, along with additional positional encoding. Extensive experiments are conducted on popular pose estimation benchmarks including Linemod, Linemod-Occlusion, and YCB-V. The results show that, VAPO improves both the keypoint correspondences and final estimated poses, and clearly achieves state-of-the-art performances.
Keywords:
6DoF pose estimation Keypoint localization Visibility
1 Introduction
Figure 1: Motivation of our visibility-aware keypoint localization method. For the input image (a), we visualize visibility of predefined keypoints in (b). Each row shows internal visibility w.r.t. self-occlusion and each column shows external visibility w.r.t. occlusion. Without direct observations in the image, localizing invisible keypoints (the yellow ones) may be challenging and degrade the quality of 3D-2D correspondences. We propose to estimate visibility-aware importance for each keypoint, and discard keypoints with low importance.
@@ 1 Introduction
Given a single input RGB image, instance-level 6DoF object pose estimator recovers rotation and translation of a rigid object with respect to a calibrated camera. The estimated pose information is crucial in many real-world applications, including robot manipulation [75, 61, 62], autonomous driving [35, 67, 29], augmented reality [36, 58], 3D reconstruction [37, 66], etc. To increase the robustness against various imaging conditions, the majority of the existing methods [46, 59, 39, 15, 44, 69, 42, 30, 51, 54, 31] first generate correspondences between 2D image pixels and 3D object points, and then regress the pose via any available Perspective-n-Point (PnP) solver [28, 64, 2].
@@
Based on the correspondence estimation process, previous methods can be divided into two categories. The first kind of methods [69, 42, 30, 64, 7, 54] estimate corresponding 3D coordinate on the object surface for each 2D pixel, which can be treated as an image-to-image translation task. The other kind of methods [46, 59, 72, 44, 31] localize predefined 3D keypoints in the input image to obtain 3D-2D correspondences. Compared with image-to-image translation based methods, keypoint-based methods efficiently encode the object geometry information, which facilitates the pose estimation process. The pioneering keypoint-based works localize the eight corners of the 3D object bounding box [46, 59]. The follow-up works [72, 44] adopt sparse keypoints (e.g., 8 keypoints) sampled from the object surface. Recently, CheckerPose [31] is proposed to generate dense correspondences via localizing dense 3D keypoints and it achieves state-of-the-art performance for object pose estimation.
@@
To obtain better correspondences for object pose estimation, a great amount of effort has been devoted to improve the localization precision of each keypoint. However, existing keypoint-based methods have a common issue illustrated in Fig. 1, that is, a large portion of predefined keypoints are invisible in the input image, due to occlusion or self-occlusion. Without direct observations, the localization results of such keypoints may be unreliable. Since the ultimate goal of keypoint localization is to establish reliable 3D-2D correspondences for 6DoF pose estimation, it may be unnecessary to localize each predefined keypoint.
@@
To overcome the above issue, we propose to estimate visibility-aware importance for each keypoint, and discard unimportant keypoints before localization. As a starting point, we can represent keypoint visibility as a binary value. We notice that such information is manually annotated in datasets from other domains, e.g., human pose estimation [21, 1]. However, annotations of keypoint visibility are currently missing in 6DoF object pose datasets. To get rid of expensive manual annotation process, we propose an efficient way to generate visibility labels from available object-level annotations. We decompose visibility into two binary terms w.r.t. external occlusion and internal self-occlusion, respectively. The external visibility term can be obtained from available object segmentation masks. The internal visibility term can be determined based on surface normals and camera ray directions, inspired by back-face culling [26, 71] in rendering. For symmetric objects, we derive modified computation to ensure consistency.
Besides binary labels, we further derive real-valued measure to comprehensively reflect the visibility-aware importance of each keypoint. To do this, we create a $k$-nearest neighbor ($k$-NN) graph from the predefined keypoints, and measure the closeness of each keypoint to visible ones as importance. We utilize Personalized PageRank (PPR) [40] as our proximity measure, and derive an analytical formula of importance which can be efficiently evaluated.
@@
Our visibility-aware importance can be easily integrated into existing keypoint-based 6DoF pose estimator to boost performance. We adopt CheckerPose [31] as our base framework, which utilizes graph neural networks (GNNs) to model the interactions among dense keypoints. While CheckerPose is originally designed to localize every keypoint in a predefined set, we add a visibility-aware importance predictor and eliminate half of the keypoints with low importance. We further enhance the keypoint embedding in GNNs with positional encoding, and use a two-stage training strategy to efficiently train the deep network.
@@
  • From object-level annotations, we derive an efficient way to generate binary keypoint visibility labels, for both asymmetric objects and symmetric ones.
  • We further derive a concise analytic formula to produce real-valued visibility-aware keypoint importance based on Personalized PageRank.
@@
  • We demonstrate that our visibility-aware importance can be easily incorporated to existing keypoint-based method to boost performance.
@@
We conduct extensive experiments on Linemod [11], Linemod-Occlusion [3], and YCB-V [68] to demonstrate the effectiveness of our method. We will release the source code of our work upon the publication.
2 Related Work
In this section, we review previous studies related to our work, including 6DoF pose estimation from RGB inputs and applications of visibility estimation.
Direct Methods. Traditionally, object poses are estimated by template matching with hand-crafted features [18, 9, 10], which can not work well on textureless objects. Features learned from neural networks are also investigated to directly produce the final estimated poses [68, 63, 53, 24, 55]. However, due to the nonlinearity of 3D rotations, direct methods are still unstable without predicting intermediate geometric representations, e.g., 3D-2D correspondences.
Image-to-image Translation Based Methods. One popular way to establish 3D-2D correspondences can be regarded as an image-to-image translation task [42]. Specifically, for each 2D pixel, the corresponding 3D point on the object surface is predicted in the object frame [42, 30, 64, 7]. The 3D coordinates can also be represented as UV-maps [69], local surface embeddings [45], hierarchical binary encodings [54], etc. The dense correspondences are robust against various imaging conditions, such as occlusions, background clutter, etc. With the correspondences, object pose can be recovered via existing PnP solvers [28, 2], coupled with RANSAC to remove outliers. Recent methods [64, 7, 33] also use neural networks to produce final poses from the correspondences.
Keypoint-based Methods. Localizing predefined keypoints in the input image is also widely used for constructing 3D-2D correspondences. For simplicity, previous works [46, 59, 15, 14, 16, 39] localize the 3D object bounding box corners. Other works [72, 44, 31] also adopt keypoints on the object surface obtained via farthest point sampling (FPS), which tend to achieve better performance since they are closely related to the object pixels. While existing methods mainly utilize sparse keypoints (e.g., 8 keypoints), CheckerPose [31] localizes dense keypoints (e.g., 512 keypoints) to construct dense 3D-2D correspondences, which increases the robustness similar to the image-to-image translation based methods. Though previous methods demonstrate the importance of predefining proper keypoints, no keypoint can always be visible in any input image.
Visibility Estimation for 3D Vision and Graphs. For an object in the 3D space, not all parts are visible in its 2D images. For rigid pose estimation, visibility can be used to select reliable correspondences [31, 17]. Still, each correspondence is generated despite its visibility. Differentiable surface visibility is developed to facilitate render-and-compare framework [49]. Visibility can also be efficiently estimated in online tracking with a known initial pose and smooth changes [27]. For nonrigid pose estimation, keypoint localization results are usually represented as heatmaps, and visibility can be implicitly interpreted from heatmaps [38, 13, 8]. Moreover, keypoint visibility is manually annotated in various datasets [21, 1, 32], and can be used as supervision signals [47, 74]. Beyond pose estimation, visibility determination also facilitates various tasks in computer vision and graphics. In rendering of large polygonal models, back-facing polygons are eliminated to speed up the rendering process [26, 71]. Visibility can also be utilized for mesh simplification [70], point cloud visualization [23, 22], novel view generation [50, 52, 56, 34, 20], multi-view aggregation [73, 6], etc.
Our work follows the keypoint-based framework for 6DoF pose estimation. We focus on keypoints with high visibility-aware importance, and effectively generate supervision signals from object-level annotations.
3 Method
@@
Our work focuses on instance-level pose estimation for a rigid object $O$ with available CAD model. We sample 3D keypoints $\mathcal{P}$ from the CAD model, and estimate the importance of each keypoint $\mathbf{p}\in\mathcal{P}$ w.r.t. visibility in the input RGB image $I$. We then localize the subset of $\mathcal{P}$ with high importance, and obtain rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ from the localization results via a PnP solver [28, 2]. We describe our method, named VAPO, in details as follows.
3.1 Generating Visibility Labels from Object-Level Annotations
Keypoint-level annotations are typically unavailable in existing 6DoF object pose datasets, such as Linemod [11], Linemod-Occlusion [3], YCB-V [68], etc. While keypoint visibility labels are provided by datasets from other domains, e.g., Leeds Sports Poses [21], MPII Human Pose [1], COCO [32], etc., such labels require expensive human labelling efforts. Moreover, there is no canonical way to predefine keypoints for 6DoF pose. Even though one can manually annotate the visibility of a specific set of keypoints, missing label issue still arises when object pose estimator utilizes a different set of keypoints, e.g., denser keypoints.
@@ 3.1 Generating Visibility Labels from Object-Level Annotations
To avoid expensive manual annotation and provide flexibility for different 3D keypoints $\mathcal{P}$, we instead seek an efficient way to generate visibility labels from available object-level annotation, i.e., rotation $\mathbf{R}$, translation $\mathbf{t}$, and object segmentation masks $M$. A keypoint $\mathbf{p}\in\mathcal{P}$ is visible if and only if it is free from both occlusions and self-occlusions. Thus we can decompose visibility into two binary terms: the external visibility $V_{\mathrm{ex}}(\mathbf{p})\in\{0,1\}$ w.r.t. occlusions from other objects, and the internal visibility $V_{\mathrm{in}}(\mathbf{p})\in\{0,1\}$ w.r.t. self-occlusions. The overall visibility $V(\mathbf{p})\in\{0,1\}$ can be computed by
@@
Since visible segmentation mask $M_{\mathrm{vis}}$ of the object $O$ reflects occlusions from other objects, we can determine $V_{\mathrm{ex}}(\mathbf{p})$ by
@@
To determine whether keypoint $\mathbf{p}\in\mathcal{P}$ is self-occluded, we can check whether the direction from $\mathbf{p}$ towards the camera has additional intersections with the object surface. However, it is time-consuming to check the intersections on the fly during training. Inspired by back-face culling [26, 71] in rendering, we compute $V_{\mathrm{in}}(\mathbf{p})$ by
$V_{\mathrm{in}}(\mathbf{p})=\begin{cases}1,&\text{if }\mathbf{d_{p}}^{(c)}\cdot\mathbf{n}_{\mathbf{p}}^{(c)}>0\\ 0,&\text{otherwise}\end{cases},$ (3)
where $\mathbf{d_{p}}^{(c)}$ denotes the direction from $\mathbf{p}$ towards the camera in the camera space, and $\mathbf{n}_{\mathbf{p}}^{(c)}$ denotes the surface normal at $\mathbf{p}$ in the camera space. In the camera space, the camera is placed at the origin $[0,0,0]^{\top}$, and then
$\mathbf{d_{p}}^{(c)}=-\mathbf{p}^{(c)}=-(\mathbf{R}\mathbf{p}+\mathbf{t}),$ (4)
where $\mathbf{p}^{(c)}$ is the keypoint coordinate in the camera space. With the 3D CAD model, we have access to the surface normal $\mathbf{n}_{\mathbf{p}}$ in the object frame, and we get
@@
thus Eq. 3 can be efficiently computed from available object-level annotations.
For a convex object, $V_{\mathrm{in}}(\mathbf{p})=1$ is a necessary and sufficient condition for $\mathbf{p}$ being not self-occluded. For a non-convex object, $V_{\mathrm{in}}(\mathbf{p})=1$ is a necessary condition, and we use Eq. 3 for several reasons. Firstly, if $V_{\mathrm{in}}(\mathbf{p})=0$, then $\mathbf{p}$ must be self-occluded, and we can safely eliminate $\mathbf{p}$ or reduce the importance of $\mathbf{p}$ in localization process. Secondly, a non-convex object tends to have a complex shape, and we might obtain better understanding of the object by keeping a small portion of self-occluded keypoints. Thirdly, we can efficiently label $V_{\mathrm{in}}(\mathbf{p})$ using available object-level annotations, which significantly facilitates the training process.
3.2 Handling Visibility of Symmetric Objects
Figure 2: Illustration of selecting the proper pose for symmetric objects. In the camera space, the camera is placed at the origin $O$ and looks at the direction $y>0$. The visible points are highlighted in red. (a) The object is translated along $y$-axis by $\mathbf{t}$. (b) The object is rotated $27^{\circ}$ counterclockwise around its center and then translated by $\mathbf{t}$. (c) The object is rotated $27^{\circ}$ clockwise and then translated by $\mathbf{t}$, but the visible points are significantly different from (b). (d) The object is rotated by larger degrees to get equivalent appearance of (c), and the visible points are more consistent with (b).
For a symmetric object, the input image corresponds to multiple equivalent poses w.r.t. the symmetry transformations $\mathcal{S}$. In practice, datasets usually provide only one annotated pose. From the perspective of keypoint visibility, however, it may not be optimal to directly use the annotated pose to generate visibility labels. Let us consider a 2D case without external occlusion for simplicity. As illustrated in Fig. 2 (a), the symmetry transformations $\mathcal{S}$ contain the identity matrix $\mathbf{I}$ and the one that rotates the object around the center for $180^{\circ}$. As shown in Fig. 2 (b), the annotated pose is a combination of a slight counterclockwise rotation around object center (denoted as $\mathbf{R}_{1}$) and a translation $\mathbf{t}$ along $y$-axis, and the corresponding visible points are on the edges $AB$ and $DA$. While in Fig. 2 (c), the annotated rotation is a slight clockwise rotation around center (denoted as $\mathbf{R}_{2}$), and the corresponding visible points appear on the edges $AB$ and $BC$. Since $|BC|=|DA|\gg|AB|$, the visible points corresponding to $(\mathbf{R}_{1},\mathbf{t})$ and $(\mathbf{R}_{2},\mathbf{t})$ are dramatically different. To make $V(\mathbf{p})$ in Fig. 2 (c) more consistent with Fig. 2 (b), we can modify the annotated pose to an equivalent one to make visible points appear on the edges $DA$ and $CD$, as shown in Fig. 2 (d).

### v5

# VAPO: Visibility-Aware Keypoint Localization for Efficient 6DoF Object Pose Estimation

@@ Abstract
Localizing predefined 3D keypoints in a 2D image is an effective way to establish 3D-2D correspondences for instance-level 6DoF object pose estimation. However, unreliable localization results of invisible keypoints degrade the quality of correspondences. In this paper, we address this issue by localizing the important keypoints in terms of visibility. Since keypoint visibility information is currently missing in the dataset collection process, we propose an efficient way to generate binary visibility labels from available object-level annotations, for keypoints of both asymmetric objects and symmetric objects. We further derive real-valued visibility-aware importance from binary labels based on the PageRank algorithm. Taking advantage of the flexibility of our visibility-aware importance, we construct VAPO (Visibility-Aware POse estimator) by integrating the visibility-aware importance with a state-of-the-art pose estimation algorithm, along with additional positional encoding. VAPO can work in both CAD-based and CAD-free settings. Extensive experiments are conducted on popular pose estimation benchmarks including Linemod, Linemod-Occlusion, and YCB-V, demonstrating that VAPO clearly achieves state-of-the-art performances. Project page: https://github.com/RuyiLian/VAPO.
I INTRODUCTION
@@
Given a single input RGB image, the instance-level 6DoF object pose estimator recovers rotation and translation of a rigid object with respect to a calibrated camera. The pose estimator is crucial in numerous real-world applications, including robot manipulation [1, 2, 3], autonomous driving [4, 5], augmented reality [6, 7], etc. To increase the robustness under various imaging conditions, most existing methods [8, 9, 10, 11, 12, 13, 14, 15, 16] first generate correspondences between 2D image pixels and 3D object points, and then regress the pose via any available Perspective-n-Point (PnP) solver [17, 18, 19].
@@
Based on the correspondence estimation process, previous methods can be divided into two categories. The first kind of methods [12, 13, 14, 18, 20, 15] estimate corresponding 3D coordinate on the object surface for each 2D pixel. The second kind of methods [8, 9, 11, 21, 16] localize predefined 3D keypoints in the input image to obtain 3D-2D correspondences, which more efficiently encode the object geometry information and facilitate the pose estimation process.
@@
To obtain better correspondences for object pose estimation, a great amount of effort has been devoted to improve the localization precision of each keypoint. However, existing keypoint-based methods share a common limitation, i.e., a large portion of predefined keypoints are invisible due to occlusion or self-occlusion. Without direct observations, localizing such keypoints often leads to unreliable results. Since the ultimate goal of keypoint localization is to establish reliable 3D-2D correspondences for 6DoF pose estimation, it may not be necessary to localize each predefined keypoint.
@@
To overcome this issue, we propose to estimate visibility-aware importance for each keypoint, and discard unimportant keypoints before localization. However, annotations of keypoint visibility are currently missing in 6DoF object pose datasets. To avoid the expensive manual annotation process, we propose an efficient way to generate visibility labels from available object-level annotations. Visibility is decomposed into two binary terms w.r.t. external occlusion and internal self-occlusion. The external visibility term can be obtained from object segmentation masks. The internal visibility term can be determined based on surface normals and camera ray directions, inspired by back-face culling [22] in rendering. For symmetric objects, we derive modified computation to ensure consistency. We further complement the binary labels with a real-valued measure which is easy to evaluate. To do this, we create a $k$-nearest neighbor ($k$-NN) graph from the predefined keypoints, and utilize Personalized PageRank (PPR) [23] to compute the closeness of each keypoint to visible ones as the desired measure.
@@
Our visibility-aware importance can be seamlessly integrated into existing keypoint-based 6DoF pose estimator to boost performance. We add a visibility-aware importance predictor before keypoint localization to eliminate keypoints with low importance. We use positional encoding to enhance the embeddings of the selected keypoints and adopt a two-stage training strategy to efficiently train our pose estimator. Furthermore, our method can easily adapt to a more general setting where precise 3D CAD model is unavailable.
@@
  • From object-level annotations, we derive an efficient way to generate binary keypoint visibility labels and real-valued visibility-aware keypoint importance, for both asymmetric objects and symmetric ones.
@@
  • We demonstrate that our visibility-aware importance can be easily incorporated to existing keypoint-based method for both CAD-based and CAD-free settings.
@@
We conduct extensive experiments on Linemod [24], Linemod-Occlusion [25], and YCB-V [26] to demonstrate the effectiveness of our method.
II Related Work
In this section, we review previous studies that are most relevant to our work. For a more comprehensive review of 6DoF object pose estimation, we refer readers to [27].
Correspondence-Based Methods for Instance-level 6DoF Pose Estimation from RGB Inputs. One popular way to establish 3D-2D correspondences can be regarded as an image-to-image translation task [13]. Specifically, for each 2D pixel, the corresponding 3D point on the object surface is predicted in the object frame [13, 14, 12, 18, 20, 15]. The dense correspondences are robust against various imaging conditions. Object pose can be recovered from the correspondences via existing PnP solvers [17, 19], coupled with RANSAC to remove outliers. Localizing predefined keypoints in the input image is also widely used for constructing 3D-2D correspondences. For simplicity, previous works [8, 9, 10] localize the 3D object bounding box corners. Other works [11, 21] also adopt sparse keypoints (e.g., 8 keypoints) on the object surface obtained via farthest point sampling (FPS). The recently proposed CheckerPose [16] localizes dense keypoints to construct dense 3D-2D correspondences, increasing the robustness similar to the image-to-image translation-based methods. 6D-Diff [28] proposes to formulate keypoint localization as a reverse diffusion process.
CAD-Free Object Pose Estimation. To remove the dependence on precise CAD models, CAD-free methods have been studied recently. RLLG [29] learns correspondences via multi-view geometric constraints. OnePose [30] obtains correspondences between the query image and SfM model via a keypoint-based matching network. OnePose++ [31] instead uses a keypoint-free feature matching pipeline for low-textured objects. GS-Pose [32] adopts 3D Gaussian Splatting [33] to build explicit representation of the object.
Visibility Estimation for 3D Vision and Graphics. Visibility has been used to select reliable correspondences for 6DoF object pose estimation [16] and relative camera pose estimation [34]. Still, each correspondence is generated despite its visibility. For nonrigid pose estimation, keypoint visibility is manually annotated in various datasets [35, 36, 37] as supervision signals. In rendering of large polygonal models, back-facing polygons are eliminated to speed up the rendering process [22].
III Method
@@
Our work focuses on instance-level pose estimation for a rigid object $O$. In the classical setting where the CAD model of $O$ is available, we sample 3D keypoints $\mathcal{P}$ over the object surface, and estimate the importance of each keypoint $\mathbf{p}\in\mathcal{P}$ w.r.t. visibility in the input RGB image $I$. We then localize the subset of $\mathcal{P}$ with high importance, and obtain rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ from the localization results via a PnP solver [17, 19]. Furthermore, our method can be easily extended to the CAD-free setting. We describe our method, named VAPO (Visibility-Aware POse estimator), in details as follows.
III-A Visibility Labels from Object-Level Annotations
Figure 1: Illustration of keypoint internal visibility. Inspired by back-face culling [22] in rendering, we compute the internal visibility based on the inner product between the directional vector towards the camera (e.g., $\mathbf{d}_{\mathbf{a}}^{(c)}$) and the normal vector (e.g., $\mathbf{n}_{\mathbf{a}}^{(c)}$).
@@ III-A Visibility Labels from Object-Level Annotations
Keypoint-level annotations are typically unavailable in existing 6DoF object pose datasets [24, 25, 26]. To avoid expensive manual annotation, we seek an efficient way to generate visibility labels from available object-level annotation, i.e., rotation $\mathbf{R}$, translation $\mathbf{t}$, and object segmentation masks $M$. A keypoint $\mathbf{p}\in\mathcal{P}$ is visible if and only if it is free from both occlusions and self-occlusions. Thus we can decompose visibility into two binary terms: the external visibility $V_{\mathrm{ex}}(\mathbf{p})\in\{0,1\}$ w.r.t. occlusions from other objects, and the internal visibility $V_{\mathrm{in}}(\mathbf{p})\in\{0,1\}$ w.r.t. self-occlusions. The overall visibility $V(\mathbf{p})\in\{0,1\}$ can be computed by
@@
Since the visible segmentation mask $M_{\mathrm{vis}}$ of the object $O$ reflects occlusions from other objects, we can determine $V_{\mathrm{ex}}(\mathbf{p})$ by
@@
To determine whether keypoint $\mathbf{p}\in\mathcal{P}$ is self-occluded, we can check whether the direction from $\mathbf{p}$ towards the camera has additional intersections with the object surface. However, it is time-consuming to check the intersections on the fly during training. Inspired by back-face culling [22] in rendering, we compute $V_{\mathrm{in}}(\mathbf{p})$ by
$V_{\mathrm{in}}(\mathbf{p})=\begin{cases}1,&\text{if }\mathbf{d}_{\mathbf{p}}^{(c)}\cdot\mathbf{n}_{\mathbf{p}}^{(c)}>0\\ 0,&\text{otherwise}\end{cases},$ (3)
where $\mathbf{d}_{\mathbf{p}}^{(c)}$ denotes the direction from $\mathbf{p}$ towards the camera in the camera space, $\mathbf{n}_{\mathbf{p}}^{(c)}$ denotes the surface normal at $\mathbf{p}$ in the camera space, and $\cdot$” indicates the dot product. As illustrated in Figure 1, keypoint $\mathbf{a}$ is internally visible while $\mathbf{b}$ is internally invisible according to Eq. (3). In the camera space, the camera is placed at the origin $[0,0,0]^{\top}$, and then
$\mathbf{d}_{\mathbf{p}}^{(c)}=-\mathbf{p}^{(c)}=-(\mathbf{R}\mathbf{p}+\mathbf{t}),$ (4)
where $\mathbf{p}^{(c)}$ is the keypoint coordinate in the camera space. When the 3D CAD model is available, we have access to the surface normal $\mathbf{n}_{\mathbf{p}}$ in the object frame, and we get
@@
thus Eq. (3) can be efficiently computed from available object-level annotations. When $V_{\mathrm{in}}(\mathbf{p})=0$, then $\mathbf{p}$ must be self-occluded, and we can safely eliminate $\mathbf{p}$ or reduce the importance of $\mathbf{p}$ in localization process. When $V_{\mathrm{in}}(\mathbf{p})=1$ and the object is non-convex, $\mathbf{d}_{\mathbf{p}}^{(c)}$ may have additional intersections with the object surface, but we omit additional inspections to accelerate the training process.
III-B Handling Visibility of Symmetric Objects
For a symmetric object, an input image corresponds to multiple equivalent poses w.r.t. the symmetry transformations $\mathcal{S}$. When the 3D CAD model is available, $\mathcal{S}$ can be obtained by manual inspection or geometric analysis. In practice, only one of the equivalent poses is annotated. For different images with similar appearances, the annotated poses may be dramatically different. From the perspective of keypoint visibility, it is not optimal to directly use the annotated poses to generate visibility labels, because the inconsistency of the labels is problematic for training a robust visibility classifier.
To enforce the consistency of visibility labels, we can transform the original annotated pose $(\mathbf{R},\mathbf{t})$ to a canonical one with a proper symmetry transformation $\tilde{\mathbf{S}}\in\mathcal{S}$. We select $\tilde{\mathbf{S}}$ by maximizing the number of internally visible keypoints in a fixed subset of keypoints $\mathcal{P}_{\mathrm{sym}}$. We consider internal visibility $V_{\mathrm{in}}$ because self-occlusion is not affected by external objects.